### v2

# Panning for Gold: Expanding Domain-Specific Knowledge Graphs with General Knowledge

Abstract.
Domain-specific knowledge graphs (DKGs) are critical yet often suffer from limited coverage compared to General Knowledge Graphs (GKGs). Existing tasks to enrich DKGs rely primarily on extracting knowledge from external unstructured data or completing KGs through internal reasoning, but the scope and quality of such integration remain limited. This highlights a critical gap: little systematic exploration has been conducted on how comprehensive, high-quality GKGs can be effectively leveraged to supplement DKGs. To address this gap, we propose a new and practical task: domain-specific knowledge graph fusion (DKGF), which aims to mine and integrate relevant facts from general knowledge graphs into domain-specific knowledge graphs to enhance their completeness and utility. Unlike previous research, this new task faces two key challenges: ① high ambiguity of domain relevance, i.e., difficulty in determining whether knowledge from a GKG is truly relevant to the target domain, and ② cross-domain knowledge granularity misalignment, i.e., GKG facts are typically abstract and coarse-grained, whereas DKGs frequently require more contextualized, fine-grained representations aligned with particular domain scenarios. To address these, we present ExeFuse, a neuro-symbolic framework based on a novel Fact-as-Program paradigm. ExeFuse treats fusion as an executable process, utilizing neuro-symbolic execution to infer logical relevance beyond surface similarity and employing target space grounding to calibrate granularity. We construct two new datasets to establish the first standardized evaluation suite for this task. Extensive experiments demonstrate that ExeFuse effectively overcomes domain barriers to achieve superior fusion performance. The source codes and datasets are available at https://github.com/eduzrh/code_datasets.
Domain-specific Knowledge Graph Fusion; Knowledge Graph Enrichment; General-to-domain Knowledge Transfer; Fact-as-Program
copyright: nonecopyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Domain-specific knowledge graphs (DKGs) have emerged as essential infrastructures for representing structured knowledge in specialized domains and enabling intelligent applications (Zhang et al., 2025; Kejriwal, 2019; Cao et al., 2024; Geisler et al., 2025; Hassani-Pak et al., 2020; Faghihi et al., 2021; Gupta et al., 2024). For example, ICEWS [^1] https://www.andybeger.com/icews captures political events for international relations analysis, while PubChem [^2] https://pubchem.ncbi.nlm.nih.gov provides chemical knowledge widely used in drug discovery. These DKGs demonstrate the great potential of domain-specific structured knowledge in powering real-world applications (Ji et al., 2022; Chen et al., 2024d; Huo et al., 2024; Zhang et al., 2024c, a). Nevertheless, compared with well-maintained general knowledge graphs (GKGs) such as Wikipedia (Erxleben et al., 2014) and YAGO (Suchanek et al., 2007), which benefit from community-driven contributions and timely updates, DKGs often suffer from limited completeness. They inherently face challenges due to restricted data sources (Lin et al., 2023), domain-specific construction methods (Khan et al., 2025; Kondylakis et al., 2025), and costly maintenance efforts (Yang et al., 2025a; Abu-Salih, 2021; Li et al., 2020; Chen et al., 2024a). These limitations create a substantial barrier for DKGs to fully support downstream applications in specialized domains (Khan et al., 2024; Paulheim, 2017). Therefore, a pressing challenge is to efficiently and accurately enrich DKGs, improving their coverage and the ability to support domain-specific applications.
Figure 1. Examples of domain-specific knowledge graph fusion. Figure 1(b) demonstrates the fusion of highly relevant entities from GKG into DKG. Figure 1(c) shows the augmented DKG.
Although currently there is no work that directly investigates how to enrich DKGs, adapting existing techniques developed for supplementing GKGs may be a viable solution. Specifically, these efforts can be broadly divided into two categories: extraction from external unstructured sources and reasoning based on internal information. ① The former category aims to acquire new knowledge from external unstructured texts, mainly including manual annotation, entity linking, and relation triple extraction (Jin et al., 2021; Nie et al., 2024; Li et al., 2024b). However, as discussed earlier, domain-specific unstructured data sources are inherently limited in scope and coverage, constraining the effectiveness of these text-based extraction methods. Moreover, the specialized terminology and complex linguistic patterns in domain-specific texts further exacerbate the challenges of accurate knowledge extraction. ② The latter category focuses on mining implicit knowledge from the existing graph itself, typically via knowledge graph completion (Wei et al., 2024), which predicts missing triples through internal inference within the existing KG. While these solutions can discover implicit relationships, the knowledge they generate is fundamentally bounded by the original KG’s scope and cannot introduce genuinely new external knowledge. Given the limitations of this existing research and the rich, continuously updated knowledge available in comprehensive GKGs like Wikipedia and YAGO, a natural question arises: Can we systematically leverage the abundant knowledge in GKGs to supplement and enrich DKGs?
To this end, we propose a new and important task setting, i.e., domain-specific knowledge graph fusion (DKGF), which aims to mine the highly relevant knowledge in the general knowledge graph to enhance the knowledge coverage of domain-specific knowledge graph, as illustrated in Figure 1. Unlike the widely studied knowledge graph alignment task (Zeng et al., 2020; Ge et al., 2021; Wu et al., 2023; Su et al., 2025; Tian et al., 2024; Ding et al., 2025), which focuses on identifying entities that are simultaneously present and semantically equivalent across different KGs, the DKGF task essentially focuses on how to extract knowledge from a large-scale semantically complex GKG that is closely related to the target domain but have not yet been included in the DKG, and then effectively integrate them, while maintaining the semantic consistency and structural integrity of the DKG itself. As a result, the DKGF task poses new challenges:
Challenge I: High Ambiguity of Domain Relevance. Firstly, GKGs contain a large amount of knowledge with low similarity to that in DKGs, yet such information may still provide critical factual associations for enriching DKGs in reality. As illustrated in Figure 1, during the mining process of relevant knowledge from GKGs, the academic institution “Brookings Institution” from Washington D.C. lacks explicit connections to entities in the political crisis event DKG (e.g., “Iran”), resulting in low semantic similarity between entities. Nonetheless, in reality, “Brookings Institution” maintains many important connections with DKG entities and is highly likely to constitute factual associations with them. This discrepancy creates ambiguity in determining whether knowledge from GKGs is truly domain-relevant, making it difficult to directly apply traditional graph matching methods or language models (Pan et al., 2024a) to this new task.
Challenge II: Cross-Domain Knowledge Granularity Misalignment. Secondly, facts associated with relevant knowledge from GKGs often appear at levels of abstraction that differ from the granularity required by DKGs, thereby making direct fusion inappropriate and even disruptive. GKGs typically encode broad, abstract facts (e.g., “[Barack Obama, head of government, United States
of America]”), whereas DKGs frequently require more contextually specific, fine-grained representations aligned with particular domain tasks (e.g., “[Barack Obama, make optimistic comment, United States of America]” in Figure 1(c)). The latter can be regarded as a contextual instantiation under the broader fact of “head of government”. Simply transferring abstract general knowledge into a DKG can introduce semantic inconsistencies and structural misalignment. Closing this granularity gap therefore requires mechanisms to transform abstract GKG facts into domain-specific representations while preserving the semantic coherence and structural integrity of the target DKG.
To tackle the root causes of DKGF, we propose ExeFuse, a neuro-symbolic framework that reframes fusion not as static matching, but as a Fact-as-Program process where knowledge facts are modeled as executable programs transformable against target domain logic. Specifically, to resolve the high ambiguity of domain relevance, ExeFuse introduces neuro-symbolic program execution, which maps discrete GKG facts into a continuous space and treats logic rules as transition operators to infer logical reachability beyond surface dissimilarity. Simultaneously, to address cross-domain knowledge granularity misalignment, we employ a target space grounding mechanism to calibrate knowledge specificity. This approach constructs a target state space and utilizes an executability core to verify that the transformed knowledge successfully grounds into the valid semantic region of the DKG, ensuring that the fused facts are not only logically relevant but also compatible with the specific granularity and structural patterns of the target domain. Notably, to fill in the gap of DKGF benchmarks, we construct two new datasets: DKGF(W-I) and DKGF(Y-I), and offer 21 benchmark configurations for systematic evaluation. Extensive experiments highlight the value of the new task and demonstrate the effectiveness of ExeFuse, providing the first standardized evaluation suite for this emerging task.
Contribution.  We summarize our contributions as follows:
  • $\bullet$
    We define a new task setting, domain-specific knowledge graph fusion (DKGF), highlighting its distinct requirements compared to traditional tasks.
  • $\bullet$
    We propose the Fact-as-Program paradigm, a neuro-symbolic approach that models fact fusion as a verifiable logical execution process in a quasi-symbolic space.
  • $\bullet$
    We provide theoretical analysis to guarantee the expressiveness of our neuro-symbolic mapping and the validity of the granularity calibration.
  • $\bullet$
    We develop two new benchmark datasets to establish a standardized evaluation for the DKGF task. Comprehensive experiments demonstrate the effectiveness, efficiency, and generalization of our ExeFuse framework.
2. Related work
To the best of our knowledge, DKGF is a novel task in domain-specific data integration. Closely related research can be categorized into extraction from external unstructured sources, internal knowledge reasoning, and cross-graph entity alignment.
Extraction from External Unstructured Sources.  This category focuses on acquiring structured knowledge from external unstructured texts to enrich KGs.
  • Relation Triple Extraction (RTE) aims to extract structured triplets from unstructured text (Hu et al., 2023; Li et al., 2024a; Yang et al., 2025b). It has evolved from supervised pipelines (Wang and Lu, 2020; Ye et al., 2022a; Lin et al., 2019) to joint extraction models. To address data scarcity in specialized domains, research has shifted towards zero-shot settings (Li et al., 2025c). Key approaches include data augmentation methods (Chia et al., 2022; Guo et al., 2024; Gong and Eldardiry, 2024) and knowledge-driven methods (Kim et al., 2023). Recently, bilevel optimization (Li et al., 2025b) and LLM-based extraction (Yang et al., 2025b) have been introduced to enhance generalization.
  • Entity Linking (EL) identifies entity mentions in text and links them to corresponding entities in KGs. While early works relied on handcrafted features (Graciotti et al., 2025), the field has moved towards dense retrieval and generative paradigms (van Hulst et al., 2020). Bi-encoder architectures (e.g., BLINK (Wu et al., 2020)) and generative EL methods (e.g., GENRE (Cao et al., 2021)) are widely used, with recent LLM-augmented pipelines (Graciotti et al., 2025; Ding et al., 2024) leveraging in-context learning.
However, both RTE and EL heavily rely on the availability and quality of domain-specific corpora, which are often scarce or noisy in specialized domains.
Internal Knowledge Reasoning.  Unlike extraction methods that seek external information, this category focuses on inferring missing facts solely based on existing information within the KG. Research in this area primarily centers on knowledge graph completion (KGC), which predicts missing triples within a single KG. Embedding-based techniques (Bordes et al., 2013; Yang et al., 2015; Trouillon et al., 2016; Sun et al., 2019; Yao et al., 2019a) project elements into continuous vector spaces to score plausibility. PLM-based methods (Yao et al., 2019a; Kim et al., 2020; Wang et al., 2021; Lv et al., 2022a; Devlin et al., 2019b) and generative approaches (Saxena et al., 2022; Chen et al., 2022a; Ye et al., 2022b) further enhance performance. Although effective for link prediction, KGC is inherently limited to the closed-world assumption of the existing graph and cannot discover new entities or facts from the outside world.
Entity Alignment.  Entity alignment (EA) aims to identify equivalent entities across KGs. Early translation-based methods (Zhao et al., 2025b; Sun et al., 2018; Bordes et al., 2013; Chen et al., 2017) utilize distance models but often struggle with complex structures. GNN-based approaches (Wang et al., 2018; Chen et al., 2022b; Mao et al., 2021; Xu et al., 2021; Cai et al., 2022; Liu et al., 2023; Zeng et al., 2024) leverage graph convolutions to capture neighborhood topologies, while recent solutions integrate side information (Wang et al., 2023a; Jiang et al., 2024b; Tang et al., 2020) or leverage LLMs for few-shot semantic matching (Chen et al., 2024b; Jiang et al., 2024a; Zhao et al., 2025a; Yang et al., 2024).
Task Distinction and Gap Analysis.  Despite these efforts, existing research predominantly relies on extracting knowledge from unstructured data or completing KGs through internal reasoning, while the scope and quality of knowledge integration remain significantly limited (Yu et al., 2022; Li et al., 2025a). This situation exposes a critical gap: There has been little systematic exploration of how comprehensive, high-quality GKGs can be effectively leveraged to supplement DKGs. DKGF is therefore an essential and indispensable task. Furthermore, this new task differs fundamentally from existing ones:
  • Vs. RTE: RTE operates on unstructured text, whereas DKGF leverages structured GKG.
  • Vs. EL: EL resolves text-to-KG disambiguation, whereas DKGF enhances DKG coverage via cross structured graph discovery and fusion.
  • Vs. KGC: KGC focuses on reasoning within a single KG, while DKGF integrates knowledge from external GKG.
  • Vs. EA: EA outputs same entity pairs across KGs, whereas DKGF identifies missing but relevant entities and produces new facts.
Consequently, existing research cannot address the core challenges of DKGF: high ambiguity of domain relevance and cross-domain granularity misalignment. While neuro-symbolic reasoning (Alshahrani et al., 2017; Cheng et al., 2024; Spillo et al., 2024; Liu et al., 2025; Cheng and Sun, 2024) provides the explicit logical grounding needed to resolve such ambiguity, it has typically been confined to single-graph inference. Inspired by this, our ExeFuse adopts a Fact-as-Program paradigm, extending neuro-symbolic execution to dynamically validate logical relevance and granularity beyond surface matching.
3. Problem Statement
In this section, we provide formal definitions for DKG and introduce the new DKGF task.
Domain-specific Knowledge Graph.  The domain-specific knowledge graphs $G^{d}=(V^{d},R^{d},F^{d})$ and general knowledge graphs $G^{g}=(V^{g},R^{g},F^{g})$ encapsulate real-world domain-specific and general knowledge through three core components: entity sets $V^{d},V^{g}$, relation sets $R^{d},R^{g}$, and fact sets $F^{d},F^{g}$ (Wang et al., 2025). Each fact in $F^{d}$ and $F^{g}$ is represented in the form of $(v^{d},r^{d},\hat{v}^{d})\in V^{d}\times R^{d}\times V^{d}$ and $(v^{g},r^{g},\hat{v}^{g})\in V^{g}\times R^{g}\times V^{g}$, respectively, where entities $v^{d},\hat{v}^{d}\in V^{d}$, $v^{g},\hat{v}^{g}\in V^{g}$, relations $r^{d}\in R^{d}$, and $r^{g}\in R^{g}$.
Domain-specific Knowledge Graph Fusion.  Based on the aforementioned definitions, the domain-specific knowledge graph fusion task aims to mine the highly relevant knowledge in the general knowledge graph $G^{g}=\big(V^{g},R^{g},F^{g}\big)$ to enhance the knowledge coverage of domain-specific knowledge graph $G^{d}=\big(V^{d},R^{d},F^{d}\big)$. Formally, given a DKG $G^{d}$ and GKG $G^{g}$, the DKGF task generates a new set of fused facts $F_{fused}$ for $G^{d}$:
(1) $F_{fused}=\{(v,r,\hat{v})\mid\;v,\hat{v}\in V^{d}\cup V^{g},\;r\in R^{d},(v,r,\hat{v})\notin F^{d}\},$
[…]
Figure 2. The framework of ExeFuse for DKGF.
4. Methodology
In this section, we present ExeFuse, a neuro-symbolic framework designed to address the unique challenges of domain-specific knowledge graph fusion (DKGF). As analyzed in Section 1, the core difficulty of DKGF lies in high ambiguity of domain relevance and cross-domain granularity misalignment. Existing static similarity matching fails to bridge this gap. Instead, effective fusion requires answering two fundamental questions: (i) how to uncover implicit logical connections hidden beneath surface-level dissimilarity, and (ii) how to adapt abstract general facts into the specific granularity of the target domain.
To this end, we propose the Fact-as-Program paradigm, which reframes fusion not as a classification task, but as a dynamic execution process (Pojer et al., 2025; Cheng et al., 2024; Alshahrani et al., 2017; Liu et al., 2025; Cheng and Sun, 2024). As illustrated in Figure 2, ExeFuse consists of three coordinated stages:
Stage 1: Latent Program Induction.  To handle structural heterogeneity, we first encode discrete GKG facts into continuous program states using a structure-aware predicate, preparing them for logical transformation.
Stage 2: Resolving Ambiguity via Neuro-Symbolic Execution.  To resolve the ambiguity of relevance, we move beyond surface embedding similarity. We treat logic rules as transition operators that “execute” the fact. By applying these operators, we infer the logical reachability of a fact—transforming it from its original state to a potential domain-relevant state. This ensures that even semantically distant entities can be linked if there exists a valid logical path.
Stage 3: Granularity Calibration via Target Space Grounding:.  To address granularity misalignment, we view the execution result as a proposal that must be verified. We construct a Target State Space representing the valid semantic distribution of the DKG. The “program” is considered successfully executed only if its output can be grounded (i.e., mapped with high probability) into this specific target region, ensuring the fused knowledge aligns with the domain’s granularity.
4.1. Formalization & Overview
In this paradigm, we treat each source GKG fact $f^{g}=(v^{g},r^{g},\hat{v}^{g})$ as a latent executable program $\Pi(f^{g})$ defined in a continuous Quasi-Symbolic Space $\mathcal{Q}\subseteq\mathbb{R}^{d}$. formally, the framework is a tuple $\Pi=(\phi_{\text{enc}},\mathcal{L},\mathcal{Q})$, where:
  • $\phi_{\text{enc}}$ is the Structure-Aware Neural Predicate that maps discrete GKG facts into continuous program states, preserving structural semantics.
  • $\mathcal{L}$ is the set of Inductive Logic Rules acting as transition operators to transform states into domain states.
  • $\mathcal{Q}$ is the differentiable state space where geometric transformations correspond to logical entailments.
A GKG fact $f^{g}$ is deemed valid for fusion into the DKG $G^{d}$ if its program is executable. Executability is defined as the ability to transition from an initial state $\mathbf{q}^{g}$ to a valid target state $\mathbf{q}^{d}$ that resides within the Target State Space of the DKG.
4.2. Latent Program Induction via Structure-Aware Predicates
The first stage compiles the discrete symbolic fact into an initial executable state. Unlike standard approaches that treat facts as flat feature vectors, we propose a Structure-Aware Neural Predicate to capture the compositional logic of the triplet $(v^{g},r^{g},\hat{v}^{g})$.
4.2.1. Structure-Aware Neural Predicate
Let $\mathbf{E}^{g}\in\mathbb{R}^{|V^{g}|\times d_{emb}}$ and $\mathbf{R}^{g}\in\mathbb{R}^{|R^{g}|\times d_{emb}}$ be the pre-trained embedding matrices. We define the initial program state $\mathbf{q}^{g}\in\mathcal{Q}$ via a relational composition function:
$\mathbf{q}^{g}=\phi_{\text{enc}}(f^{g})=\mathbf{W}_{2}\cdot\text{GeLU}\left(\mathbf{W}_{1}\left[(\mathbf{e}_{v^{g}}\odot\mathbf{e}_{r^{g}})\,;\,(\mathbf{e}_{\hat{v}^{g}}\odot\mathbf{e}_{r^{g}})\,;\,\mathbf{e}_{r^{g}}\right]+\mathbf{b}_{1}\right),$
where $\mathbf{e}_{(\cdot)}$ denotes the embedding vector retrieved from $\mathbf{E}^{g}$ or $\mathbf{R}^{g}$, $\odot$ represents the Hadamard (element-wise) product, and $[;]$ denotes concatenation. $\mathbf{W}_{1},\mathbf{W}_{2}$ and $\mathbf{b}_{1}$ are learnable parameters. Remark: The interaction terms $(\mathbf{e}_{v}\odot\mathbf{e}_{r})$ and $(\mathbf{e}_{\hat{v}}\odot\mathbf{e}_{r})$ explicitly model the interaction between the relation and the head/tail entities, ensuring that $\mathbf{q}^{g}$ encodes the structural validity of the fact.
4.2.2. Target State Space
We define the Target State Space $\mathcal{Q}^{d}\subset\mathcal{Q}$ as the semantic region populated by the embeddings of all existing valid DKG facts:
(3) $\mathcal{Q}^{d}=\{\phi_{\text{dkg}}(f^{d})\mid f^{d}\in F^{d}\},$
where $\phi_{\text{dkg}}$ is the domain-specific encoder. It shares the same architecture as $\phi_{\text{enc}}$ but is fine-tuned to represent the target domain distribution.
4.3. Stage 2: Resolving Ambiguity via Neuro-Symbolic Execution
This stage addresses Challenge I (Ambiguity of Domain Relevance). In a static space, relevance is often ambiguous because direct similarity fails to capture implicit connections. We argue that relevance is best established through logical reachability (Spillo et al., 2024): if a fact implies a valid domain state through explicit reasoning rules, the ambiguity is resolved.
4.3.1. Differentiable Logic Rules
We learn a set of logic rules $\mathcal{L}$. Symbolically, a rule $R\in\mathcal{L}$ implies $r^{d}(X,Y)\leftarrow r^{g}(X,Y)\wedge\mathcal{C}$. To support differentiable execution, we parameterize each rule $R$ as an affine transformation. The rule application function $\mathcal{T}_{R}:\mathcal{Q}\to\mathcal{Q}$ is defined as:
(4) $\mathcal{T}_{R}(\mathbf{q}^{g})=\mathbf{w}_{R}\odot\mathbf{q}^{g}+\mathbf{b}_{R},$
where $\mathbf{w}_{R},\mathbf{b}_{R}\in\mathbb{R}^{d}$ are learnable parameters. Geometrically, this transformation models the inference step, where $\mathbf{w}_{R}$ and $\mathbf{b}_{R}$ adjust the semantic focus to trace the logical path from the source fact to the target context.
4.3.2. Neural Rule Selector
To address the ambiguity of which rule applies, we employ a Neural Rule Selector that computes an attention distribution $p(R\mid\mathbf{q}^{g})$ over $\mathcal{L}$:
(5) $p(R\mid\mathbf{q}^{g})=\frac{\exp(\mathbf{u}_{R}^{\top}\tanh(\mathbf{W}_{sel}\mathbf{q}^{g}))}{\sum_{R^{\prime}\in\mathcal{L}}\exp(\mathbf{u}_{R^{\prime}}^{\top}\tanh(\mathbf{W}_{sel}\mathbf{q}^{g}))},$
where $\mathbf{W}_{sel}\in\mathbb{R}^{d\times d}$ projects the state into the rule-selection space, and $\mathbf{u}_{R}\in\mathbb{R}^{d}$ is a rule-specific query vector.
4.3.3. Program Execution
The executed program state $\hat{\mathbf{q}}^{d}$ is the expected outcome of applying the rules:
(6) $\hat{\mathbf{q}}^{d}=\sum_{R\in\mathcal{L}}p(R\mid\mathbf{q}^{g})\cdot\mathcal{T}_{R}(\mathbf{q}^{g}).$
By explicitly executing these logic rules, ExeFuse bridges the ambiguity gap via logical inference, recovering implicit connections that pure similarity search would miss.
4.4. Stage 3: Granularity Calibration via Target Space Grounding
This stage addresses Challenge II (Granularity Misalignment). While logic execution ensures relevance, the resulting state $\hat{\mathbf{q}}^{d}$ may still be too abstract (e.g., generic concepts) to fit the DKG. To solve this, we perform Granularity Calibration by checking if the executed state $\hat{\mathbf{q}}^{d}$ can be successfully grounded into the Target State Space $\mathcal{Q}^{d}$.
4.4.1. Executability Score
We define the executability score $S_{exec}(f^{g})$ using kernel density estimation (KDE) over a set of DKG prototypes $F^{d}_{sub}\subset F^{d}$:
(7) $S_{exec}(f^{g})=\frac{1}{|F^{d}_{sub}|}\sum_{f^{d}_{j}\in F^{d}_{sub}}\exp\left(-\frac{\|\hat{\mathbf{q}}^{d}-\phi_{\text{dkg}}(f^{d}_{j})\|^{2}}{\tau}\right),$
where $\tau$ is a temperature hyperparameter. To ensure efficiency, $F^{d}_{sub}$ consists of $K$ cluster centroids derived from K-Means clustering on the DKG fact embeddings ($K\ll|F^{d}|$). Intuitively, $S_{exec}$ measures whether the transformed fact “lands” in a semantically valid region of the Target State Space.
4.4.2. Fusion Decision
The final fusion probability is computed by a fusion network $\Psi$ that integrates the structural context and the verification score:
(8) $p(\text{fuse}\mid f^{g})=\sigma\left(\Psi([\mathbf{q}^{g}\,;\,\hat{\mathbf{q}}^{d}\,;\,\tilde{S}_{exec}])\right),$
where $\tilde{S}_{exec}$ is the normalized executability score. Facts with $p>\delta$ are added to the final fused set $F_{fused}$.
4.5. Training Objective
We optimize the framework using a joint loss function on a dataset of aligned pairs $\mathcal{D}_{pos}$ (constructed via distant supervision) and negative samples.
Main Fusion Loss. We minimize the binary cross-entropy loss $\mathcal{J}_{main}$ for the fusion decision $p(\text{fuse})$:
(9) $\mathcal{J}_{main}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})\right].$
Rule Consistency Loss. To enforce valid semantic transformations, we minimize the regression loss on positive pairs:
(10) $\mathcal{J}_{rule}=\sum_{(f^{g},f^{d})\in\mathcal{D}_{pos}}\|\hat{\mathbf{q}}^{d}(f^{g})-\phi_{\text{dkg}}(f^{d})\|^{2}.$
The final objective is $\mathcal{J}=\mathcal{J}_{main}+\lambda\mathcal{J}_{rule}$.
4.6. Complexity Analysis
We analyze the inference complexity to demonstrate the scalability of ExeFuse. Let $|F^{g}|$ be the number of GKG facts, $d_{emb}$ be the embedding dimension, and $K$ be the number of prototypes. The complexity per fact consists of:
  • Latent Program Induction: $O(d_{emb}^{2})$ for the structure-aware neural predicate, dominated by matrix-vector multiplications.
  • Neuro-Symbolic Execution: $O(|\mathcal{L}|\cdot d_{emb})$ for attention calculation and element-wise affine transformations. Since $|\mathcal{L}|$ is small (typically <100<100), this step is highly efficient.
  • Target Space Grounding: $O(K\cdot d_{emb})$ for measuring the distance against $K$ prototypes in the Target State Space. Using clustered prototypes ($K\ll|F^{d}|$) avoids the prohibitive cost of traversing the entire DKG.
Consequently, the overall complexity is $O(|F^{g}|\cdot(d_{emb}^{2}+|\mathcal{L}|d_{emb}+Kd_{emb}))$. This scales linearly with the GKG size $|F^{g}|$, offering a significant efficiency advantage over traditional pair-wise matching methods that typically require quadratic $O(|F^{g}|\cdot|F^{d}|)$ operations.
4.7. Theoretical Motivation
To justify the design choices of ExeFuse, we analyze the expressive power of the neuro-symbolic mapping and the geometric interpretation of the grounding mechanism.
Proposition 0 (Expressive Completeness of Neuro-Symbolic Mapping).
Let $\mathcal{T}:\mathcal{Q}\to\mathcal{Q}$ be the ideal transition function mapping source facts to the target domain. The affine rule operators $\mathcal{L}$, combined with the non-linear neural selector $\mathcal{P}$, constitute a universal approximator. Specifically, the expected execution $\hat{q}^{d}=\sum_{R\in\mathcal{L}}p(R|q^{g})\cdot(w_{R}\odot q^{g}+b_{R})$ can approximate any continuous function on a compact set to arbitrary precision, given sufficient rules $|\mathcal{L}|$.
Analysis. Proposition 1 justifies the use of multiple discrete logic rules instead of a single projection matrix. While a single affine transformation is limited to linear mappings, our mechanism functions as a mixture of affine experts. The neural selector dynamically weights these experts, enabling the model to fit complex, non-linear decision boundaries required to bridge the semantic gap between GKG and DKG, ensuring that valid domain states are reachable within the model’s hypothesis space.
Proposition 0 (Granularity Alignment via Density Maximization).
Let $p_{\text{data}}(x)$ be the probability density function of valid facts in the target DKG space $\mathcal{Q}^{d}$. The executability score $S_{exec}(f^{g})$, defined via kernel density estimation over prototypes, acts as a proxy for the log-likelihood of the fused fact under the target distribution. Maximizing $S_{exec}(f^{g})$ is equivalent to minimizing the divergence between the generated fact embedding and the underlying data manifold of $G^{d}$.
Analysis. Proposition 2 clarifies that $S_{exec}$ is not merely a distance heuristic but a density-based verification mechanism. By optimizing this score, ExeFuse explicitly constrains the fused facts to reside in high-density regions of the target space (i.e., the valid semantic manifold of the domain). This theoretically guarantees that the generated facts are not just logically relevant but also conform to the specific granularity and structural patterns characteristic of the DKG.
Table 1. Dataset statistics. $F_{fused}$”: The total number of fused new facts (tuples) from GKG into DKG. “Structure. Sim.”: The average neighbor structure similarity of entities in DKG and GKG, as defined in (Jiang et al., 2024b). “#Rel. O.”: The proportion of overlapping relations in DKG and GKG.
[…]
Table 2. Main experiment results on DKGF(W-I) and DKGF(Y-I) datasets. We report the accuracy (ACC), precision (P), recall (R), and F1-score (F1) for each method. Since ACC and F1 provide a comprehensive evaluation of model performance (Zhang et al., 2024d), the best ACC/F1 results are highlighted in bold, while the runner-up results are underlined. “ Sema., Struc., LLM.” indicate the use of semantics information, structural information, and large language models, respectively.
Benchmark Configurations Settings DKGF(W-I)-S1 DKGF(Y-I)-S1 DKGF(W-I)-S2 DKGF(Y-I)-S2 Sema. Struc. LLM. ACC P R F1 ACC P R F1 ACC P R F1 ACC P R F1 General-purpose Rule. StringMatch-F 0.496 0.160 0.002 0.004 0.498 0.286 0.003 0.005 0.498 0.192 0.002 0.003 0.498 0.308 0.003 0.005 TF-IDF-F 0.498 0.461 0.025 0.047 0.507 0.598 0.043 0.081 0.498 0.451 0.023 0.043 0.508 0.598 0.046 0.086 Trans. TransE-F 0.649 0.668 0.596 0.630 0.588 0.782 0.244 0.372 0.637 0.662 0.560 0.607 0.575 0.772 0.212 0.332 TransH-F 0.640 0.683 0.520 0.591 0.583 0.764 0.240 0.365 0.632 0.685 0.489 0.570 0.568 0.768 0.195 0.311 DistMult-F 0.554 0.547 0.629 0.585 0.525 0.524 0.548 0.536 0.536 0.532 0.590 0.560 0.532 0.530 0.562 0.545 ComplEx-F 0.553 0.546 0.632 0.586 0.526 0.525 0.567 0.545 0.544 0.539 0.609 0.572 0.515 0.514 0.546 0.530 GNN. GCN-F 0.485 0.487 0.572 0.526 0.501 0.501 0.793 0.614 0.493 0.494 0.582 0.534 0.498 0.499 0.796 0.613 TransGNN-F 0.490 0.491 0.505 0.498 0.465 0.465 0.477 0.471 0.486 0.487 0.507 0.497 0.468 0.468 0.481 0.474 Graph-Memba-F 0.485 0.487 0.572 0.526 0.503 0.502 0.794 0.615 0.493 0.494 0.582 0.534 0.499 0.501 0.796 0.615 Generative. BERT-F 0.532 0.531 0.549 0.540 0.586 0.599 0.523 0.558 0.530 0.531 0.507 0.519 0.569 0.561 0.631 0.594 ICL-F 0.550 0.553 0.528 0.540 0.583 0.573 0.649 0.609 0.494 0.494 0.523 0.508 0.488 0.489 0.523 0.505 Self-Consistency-F 0.592 0.607 0.522 0.561 0.590 0.584 0.621 0.602 0.531 0.533 0.507 0.520 0.566 0.559 0.628 0.591 Self-RAG-F 0.576 0.589 0.506 0.544 0.594 0.623 0.477 0.540 0.544 0.544 0.545 0.545 0.577 0.595 0.481 0.532 Cross-task Adaptation EA. SimpleHHEA-F 0.490 0.492 0.507 0.499 0.493 0.493 0.503 0.498 0.492 0.492 0.507 0.500 0.477 0.477 0.481 0.479 ChatEA-F 0.596 0.551 0.581 0.566 0.649 0.727 0.477 0.576 0.592 0.611 0.507 0.554 0.592 0.619 0.481 0.541 KGC. KG-BERT-F 0.523 0.522 0.546 0.534 0.516 0.515 0.531 0.523 0.507 0.507 0.515 0.511 0.503 0.503 0.481 0.492 KG-LLaMA-F 0.512 0.512 0.531 0.521 0.539 0.537 0.558 0.547 0.529 0.527 0.529 0.528 0.521 0.521 0.514 0.517 KoPA-F 0.558 0.556 0.577 0.566 0.561 0.556 0.602 0.578 0.549 0.547 0.564 0.556 0.541 0.543 0.522 0.532 PRGC-F 0.503 0.503 0.508 0.505 0.507 0.507 0.515 0.511 0.504 0.504 0.507 0.506 0.503 0.503 0.481 0.492 RTE. NoGen-BART-F 0.516 0.515 0.531 0.523 0.511 0.511 0.523 0.517 0.508 0.508 0.517 0.512 0.509 0.509 0.491 0.500 NoGen-T5-F 0.510 0.510 0.522 0.516 0.507 0.507 0.515 0.511 0.509 0.509 0.519 0.514 0.504 0.504 0.509 0.506 ExeFuse (Ours) 0.680 0.673 0.708 0.690 0.661 0.615 0.717 0.662 0.655 0.627 0.683 0.654 0.633 0.634 0.682 0.657
Table 3. Transferability analysis on DKGF(W-I)-S1.
Model Seen Unseen All
ACC F1 ACC F1 ACC F1
DistMult-F 0.563 0.623 0.491 0.340 0.554 0.585
Self-RAG-F 0.565 0.570 0.594 0.372 0.576 0.544
ChatEA-F 0.615 0.610 0.675 0.429 0.632 0.589
ExeFuse (Ours) 0.672 0.709 0.731 0.565 0.680 0.690
5. Experiments
In this section, we introduce the experimental setup and then extensively discuss the results, highlighting the strengths of our proposed datasets and framework from multiple perspectives. The following six research questions (RQ) are the focus of our experiments.
  • RQ1.
    Can ExeFuse effectively address the challenges of challenges and achieve significant improvements?
  • RQ2.
    Does the new task offer sufficient research value, and do the DKGF(W-I) and DKGF(Y-I) datasets provide meaningful insights for the solution to this new task?
  • RQ3.
    How much do each module in the ExeFuse contribute to the final results? Are these modules reasonably designed?
  • RQ4.
    Does the ExeFuse successfully balance alignment accuracy and efficiency?
  • RQ5.
    Which aspects of this challenging task require further in-depth investigation?
  • RQ6.
    Are there intuitive cases to straightly demonstrate the effectiveness of ExeFuse?
5.1. Experimental Settings
5.1.1. DKGF Benchmarks
Currently, there is no benchmark that effectively studies and evaluates the DKGF task. To this end, we construct two new DKGF benchmarks, i.e., DKGF(W-I) and DKGF(Y-I), designed to support knowledge integration and application across domains. As shown in Table 1, the DKG is drawn from the Integrated Crisis Early Warning System (ICEWS) database, which provides data on temporal political events that capture evolving geopolitical relationships among political actors. General knowledge comes from Wikidata and YAGO, both of which are high-quality, large-coverage resources that provide rich background facts and entity attributes. During dataset construction, we align and fuse cross-domain knowledge by incorporating relevant information from Wikidata and YAGO into ICEWS. This process mitigates the incompleteness of domain knowledge, improves multidimensional event understanding, and offers more comprehensive data support for knowledge reasoning and applications in integrated crisis early warning tasks.
5.1.2. Benchmark Suite Construction & Adaptation
Since DKGF is a nascent task with no pre-existing specialized solutions, a primary contribution of this work is to establish a comprehensive and standardized evaluation suite to facilitate future research. We systematically adapted 21 representative methods from related fields (e.g., KGC, Entity Alignment) to the DKGF setting using a unified Fusion Scoring Protocol. Specifically, to bridge the gap between their original tasks and DKGF: (1) for translation and GNN-based methods (e.g., TransE, GCN), we retrained them on the joint graph $G^{d}\cup G^{g}$ to score candidate fusion triples $(h^{d},r^{d},t^{g})$ based on link plausibility; (2) for Entity Alignment methods (e.g., ChatEA), we utilized aligned entity pairs to project potential facts from $G^{g}$ into the $G^{d}$ semantic space. This rigorous adaptation not only ensures a fair comparison but also provides a robust and reproducible benchmark configuration ecosystem for the community. These configurations are categorized as follows:
  • $\bullet$
    General-purpose Configurations. To establish benchmark performance using universal architectures, this category employs broadly applicable methods (e.g., TransE, GNN, BERT, LLMs) rather than task-specific designs. This category of configurations is mainly based on current advanced and classic general methods, including rule-based methods (i.e., “Rule.”), such as StringMatch-F (Du et al., 2025) and TF-IDF-F (Paulsen et al., 2023); Translation-based methods (i.e., “Trans.”) such as TransE-F (Bordes et al., 2013), TransH-F (Sun et al., 2019), DistMult-F (Yang et al., 2015), and ComplEx-F (Trouillon et al., 2016); GNN-based methods (i.e., “GNN”) like GCN-F (Chen et al., 2024c; Zhang et al., 2022), TransGNN-F (Zhang et al., 2024b), and Graph-Memba-F (Behrouz and Hashemi, 2024); Generative methods (i.e., “Generative.”) like BERT-F (Devlin et al., 2019a), ICL-F (Kothapalli et al., 2025; Wei et al., 2022), Self-Consistency-F (Wang et al., 2023b; Pan et al., 2024b), and Self-RAG-F (Asai et al., 2024);
  • $\bullet$
    Cross-task Adaptation Configurations. This category of configuration mainly involves improving representative methods from current related research tasks to adapt to the DKGF task, including entity alignment (i.e., “EA.”), such as SimpleHHEA-F (Jiang et al., 2024b) and ChatEA-F (Jiang et al., 2024a); knowledge graph completion (i.e., “KGC.”), such as KG-BERT-F (Yao et al., 2019b; Zhang et al., 2024d), KG-LLaMA-F (Yao et al., 2025; Zhang et al., 2024d), KoPA-F (Zhang et al., 2024d) and PRGC-F (Lv et al., 2022b); and relation triple extraction (i.e., “RTE.”), like NoGen-BART-F (Li et al., 2024b), and NoGen-T5-F (Li et al., 2024b).
Baseline adaptation protocol.
Adapting existing methods to DKGF is non-trivial, as the task involves cross-graph fusion without shared embedding spaces. For each baseline, we follow its strongest feasible adaptation under a unified information setting, without introducing task-specific heuristics unavailable to other methods. For example, translation-based KGC models (e.g., TransE) are trained on the unified graph $G^{d}\cup G^{g}$ with disjoint entity embeddings, and candidate triples $(h^{d},r^{d},t^{g})$ are scored using their standard plausibility functions. Entity alignment methods leverage aligned entity pairs to project candidate facts, while RTE-based methods operate purely on structured triples without access to gold fusion labels.
5.1.3. Evaluation Protocol and Implementation Details
To evaluate the quality of DKGF produced by ExeFuse and the benchmark configurations, we follow the triple classification task (Li et al., 2024b; Zhang et al., 2024d) and use accuracy (ACC), precision (P), recall (R), and F1-score (F1) as evaluation metrics. This task is essentially a binary classification problem, and all test datasets are label-balanced. We also follow two classic data split settings (Zhang et al., 2024d; Jiang et al., 2024a), where DKGF(W-I)-S1 and DKGF(Y-I)-S1 use 80% of training data, and DKGF(W-I)-S2 and DKGF(Y-I)-S2 use 70% to test robustness. We use runtime (in seconds) as an efficiency metric. All LLMs are implemented using the same model version, GPT-4 (gpt-4-0125-preview).
5.2. Main Results (RQ1 & RQ2)
5.2.1. Comparison with All Configurations
As shown in Table 2, we have performed a comprehensive comparison for this new task. These results lead to several key observations:
  • $\bullet$
    Comparison with the General-purpose Configurations. We first compare ExeFuse with general-purpose methods, including rule-based, translation-based, GNN-based, and generative configurations. These methods perform substantially worse than ExeFuse across all settings. This indicates that directly modeling semantic similarity or structural proximity is insufficient for domain-specific knowledge graph fusion, where relevance cannot be reliably inferred without considering domain-dependent execution constraints.
  • $\bullet$
    Comparison with the Cross-task Adaptation Configurations. We also compare ExeFuse with methods adapted from entity alignment, knowledge graph completion, and relation triple extraction tasks. Despite their effectiveness in their original tasks, these methods fail to achieve competitive performance and ExeFuse consistently achieves state-of-the-art performance in DKGF. This suggests that DKGF is not a straightforward extension of existing KG-related tasks, and that directly transferring models without task-specific relevance modeling is inadequate.
  • $\bullet$
    Overall Observation. Across all configurations, ExeFuse achieves up to 9.5% relative improvement in ACC and F1 over the strongest benchmark configurations. These results demonstrate that DKGF requires explicitly modeling domain relevance and cross-domain abstraction gaps, which are not adequately addressed by existing general-purpose or cross-task adaptation configurations.
  • $\bullet$
    Evaluation of Task Value and Dataset Challenge. Furthermore, we perform a comprehensive analysis of the DKGF(W-I) and DKGF(Y-I) datasets, focusing on their complexity of fusion and practical utility. As shown in Table 2, the results indicate a significant potential for improvement in overall performance, as evidenced by the F1 scores of the state-of-the-art benchmark configurations, which are only 0.607 and 0.615 for DKGF(W-I)-S2 and DKGF(Y-I)-S2, respectively. It is worth noting that the performance of most models in related research tasks, such as entity alignment and knowledge graph completion, exceeds 0.9 in many cases (Jiang et al., 2024a; Zhang et al., 2024d). These findings highlight the inherent difficulty of the task and suggest that there are still numerous unresolved challenges.
5.2.2. Transferability Exploration
To verify that ExeFuse learns transferable logic rather than memorizing entities, we evaluated performance on unseen entities. Specifically, we partition the test set facts into two categories: seen and unseen. Specifically, the facts in the test set are divided into two parts: the seen part and the unseen part. If the head or tail entity in a fact is not present in the training set, it is regarded as unseen. In this setting, a subset of entities does not participate in the training process, making it difficult for the model to locate these truly relevant entities, which makes the test process more challenging. As shown in Table 3, ExeFuse significantly outperforms baselines on unseen data, maintaining consistently high F1 scores. In contrast, methods like ChatEA-F exhibit instability when facing entities not present during training. This confirms that our neuro-symbolic execution paradigm generalizes to new geometric regions of the DKG.
Table 4. Ablation study and design verification on DKGF(W-I)-S1. “Avg.”: average of ACC and F1. $\Delta$”: relative performance drop in Avg.
Variant DKGF(W-I)-S1 $\Delta$
ACC F1 Avg.
ExeFuse (Full Model) 0.680 0.690 0.685 -
I. Complexity Validation (Is the complex pipeline necessary?)
w/o Rules $\mathcal{L}$ (Single Affine Projection) 0.605 0.612 0.609 -11.1%
w/o Reasoning (Direct Classification) 0.577 0.583 0.580 -15.3%
II. Module Necessity (Do all components contribute?)
w/o Structure-Aware Encoder ($\phi_{enc}$) 0.570 0.582 0.576 -15.9%
w/o Neural Rule Selector ($p(R|q^{g})$) 0.620 0.623 0.622 -9.2%
w/o Granularity Calibration ($S_{exec}$) 0.595 0.601 0.598 -12.7%
III. Design Choice Verification (Are choices optimal?)
Replace Affine Rules with MLP 0.665 0.672 0.669 -2.3%
Replace KDE with Min-Distance 0.648 0.655 0.652 -4.8%
Table 5. Accuracy and cost analysis of ExeFuse and benchmark configurations on DKGF(W-I)-S1. Avg.tokens/Avg.time: the average tokens/time cost of the model.
Settings DKGF(W-I)-S1 (Accuracy) Cost (Efficiency and Effectiveness)
ACC F1 Avg. Avg. tokens Avg. time (seconds)
Self-Consistency-F 0.592 0.561 0.577 41,561,586 178,041.6
ChatEA-F 0.596 0.566 0.581 48,162,276 254,417.4
TransH-F 0.640 0.591 0.583 0 55.7
SimpleHHEA-F 0.490 0.499 0.495 0 2628.2
ExeFuse (Ours) 0.680 0.690 0.685 0 42.6
Table 6. Performance analysis of the new task “Relevant Entity Finding (REF)”.
Models DKGF(W-I)-S1 DKGF(Y-I)-S1 DKGF(W-I)-S2 DKGF(Y-I)-S2
ACC P R F1 Avg. ACC P R F1 Avg. ACC P R F1 Avg. ACC P R F1 Avg.
DistMult-F 0.975 0.830 0.840 0.835 0.905 0.946 0.853 0.730 0.787 0.866 0.974 0.845 0.869 0.857 0.916 0.936 0.863 0.772 0.815 0.876
Self-RAG-F 0.970 0.849 0.731 0.785 0.878 0.936 0.897 0.603 0.721 0.829 0.969 0.833 0.819 0.826 0.898 0.921 0.885 0.649 0.749 0.835
ChatEA-F 0.973 0.888 0.737 0.806 0.889 0.938 0.937 0.583 0.719 0.828 0.969 0.868 0.769 0.815 0.892 0.921 0.895 0.642 0.747 0.834
ExeFuse 0.982 0.968 0.791 0.871 0.927 0.958 0.958 0.721 0.823 0.890 0.978 0.960 0.787 0.865 0.922 0.955 0.896 0.852 0.873 0.914
5.3. Ablation Study (RQ3)
To rigorously evaluate the necessity of our architecture, we conducted a comprehensive ablation study on DKGF(W-I)-S1 (see Table 4). Specifically, we verify how each component contributes to addressing the two core DKGF challenges.
5.3.1. Justification of Complexity vs. Performance
We first address the concern of whether the neuro-symbolic complexity is warranted.
  • Single Affine Projection: We replaced the multi-rule logic execution with a single learnable linear transformation matrix. The performance drops significantly ($\Delta$-11.1%) to an F1 of 0.612. Notably, this result is comparable to the translation-based baseline TransE-F (F1 0.630) in Table 2. This confirms that simple linear mappings hit a performance bottleneck and are insufficient to bridge the complex semantic gap between GKG and DKG.
  • Direct Classification: Removing the reasoning module entirely leads to the worst performance ($\Delta$-15.3%), proving that the encoder alone cannot resolve the ambiguity of relevance without explicit logical transformation.
5.3.2. Impact of Key Modules & Challenge Resolution
We validate the contributions of each stage in the Fact-as-Program paradigm and their roles in solving specific tasks:
  • Latent Program Induction: Replacing the Structure-Aware Predicate ($\phi_{enc}$) with a standard linear encoder leads to the largest drop ($\Delta$-15.9%). This confirms that treating facts as structured programs, preserving the compositional logic of $(h,r,t)$, is the foundational basis for executable reasoning.
  • Neuro-Symbolic Execution (Addressing Challenge I): Removing the dynamic rule selector (using uniform weights) degrades performance by 9.2%, while removing execution entirely causes a 15.3% drop. This demonstrates that static similarity matching fails to capture implicit logical connections. The proposed neuro-symbolic execution effectively resolves Challenge I (Ambiguity of Domain Relevance) by inferring logical reachability beyond surface text.
  • Target Space Grounding (Addressing Challenge II): Removing the granularity calibration mechanism ($S_{exec}$) leads to a 12.7% drop. This finding is critical: it indicates that even if a fact is logically relevant, it may still fail to fuse correctly due to Challenge II (Granularity Misalignment). The grounding mechanism acts as a necessary compiler check, ensuring the transformed knowledge aligns with the specific density and granularity of the target domain.
5.3.3. Verification of Design Choices
Finally, we justify specific neuro-symbolic reasoning technical decisions:
  • Affine vs. MLP Rules: Replacing affine operators with non-linear MLPs degrades performance by 2.3%. We attribute this to the higher parameter count of MLPs leading to overfitting on the sparse DKG data, whereas affine transformations offer a robust and parameter-efficient representation for logic rules.
  • KDE vs. Min-Distance: Replacing KDE-based density estimation with simple Euclidean distance to the nearest centroid causes a 4.8% drop. This confirms that density-based verification is more reliable for quantifying granularity alignment than point-to-point distance metrics.
5.4. Efficiency Analysis (RQ4)
We evaluate ExeFuse’s efficiency against two categories: ① lightweight general-purpose and cross-task adaptation configurations (e.g., TransH-F, SimpleHHEA-F), which are efficient but limited in capturing complex domain interactions, and ② LLM-based general-purpose and cross-task adaptation configurations (e.g., ChatEA-F, Self-Consistency-F), which exploit LLMs but incur high computational overhead. As shown in Table 5 , ExeFuse achieves the best trade-off between effectiveness and efficiency: it delivers the highest accuracy (Avg. 0.685) while eliminating token consumption through our small-model paradigm, requiring only 42.6 seconds with substantial performance gains (17.9%-38.4%). This validates the advantage of our Fact-as-Program paradigm, which provides a new perspective for DKGF by reformulating knowledge fusion as executable neuro-symbolic reasoning without relying on expensive LLM API calls. These results validate its robustness in overcoming performance bottlenecks on the DKGF benchmark. The further analysis of theoretical complexity can be found in Section 4.6.
5.5. Relevant Entity Finding (RQ5)
In the DKGF process, accurate identification of general knowledge relevant to the DKG is a critical prerequisite to achieve reliable fusion. Failure to discover the correct cross-domain related entities results in erroneous fact generation and downstream inference. To address RQ5, we formulate and systematically study an essential yet under-explored task in the DKGF pipeline, namely Relevant Entity Finding (REF), which aims to identify a set of entities in the GKG that are semantically related to the DKG. We compare ExeFuse against embedding-based (DistMult-F) and LLM-based (Self-RAG-F, ChatEA-F) benchmark configurations. As shown in Table 6, we observe: (1) Resolving Ambiguity via Logical Reachability: ExeFuse significantly outperforms ChatEA-F (0.871 vs. 0.806 F1 on DKGF(W-I)-S1). While embeddings rely on surface similarity and LLMs lack structural constraints, ExeFuse redefines relevance as Logical Reachability. An entity is deemed relevant only if its associated fact successfully transforms into a valid DKG target state via neuro-symbolic execution. (2) Correlation with Fusion: Performance trends in REF closely mirror the main DKGF results in Table 6. This correlation validates our hypothesis that resolving domain relevance ambiguity (Stage 2) is the fundamental prerequisite for effective knowledge fusion.
Figure 3. The venn diagram of correct predictions on DKGF(W-I)-S1. Intersections show shared same predictions.
5.6. Case study (RQ6)
To intuitively demonstrate ExeFuse’s effectiveness, we conduct a multi-level analysis combining macro-statistics and micro-cases. (1) Macro Analysis: The Venn diagram in Figure 3 shows that ExeFuse correctly identifies a unique subset of fused facts ($27.9\%$) missed by both embedding-based and LLM-based baselines. This superiority stems from our neuro-symbolic approach, which captures complex logical dependencies and granularity alignments inaccessible to methods relying solely on surface text or static embeddings. (2) Micro Analysis: Consider the challenging inference: (Turki bin Faisal Al Saud, Host a visit, Brookings Institution). While DistMult-F fails due to low surface similarity, and LLMs (e.g., ChatEA-F) fail to bridge the semantic gap between the retrieved GKG fact (employer) and the target DKG relation (Host a visit), ExeFuse succeeds by treating the GKG fact as an executable program. Our Neuro-Symbolic Execution identifies the latent logic bridging these relations, while Target Space Grounding ensures the result aligns with the fine-grained granularity of political events. This confirms ExeFuse’s ability to resolve ambiguity through logical reachability.
6. Conclusion and Future Work
This paper introduces domain-Specific knowledge graph fusion (DKGF), a novel task aimed at enriching specialized knowledge graphs with general knowledge while overcoming the challenges of relevance ambiguity and granularity misalignment. To address these issues, we propose ExeFuse, a neuro-symbolic framework based on the Fact-as-Program paradigm. ExeFuse treats knowledge fusion as an executable logical process, employing neuro-symbolic execution to infer deep relevance and a target space grounding mechanism to calibrate knowledge granularity. We also establish the first standardized evaluation suite with two new benchmark datasets. Extensive experiments demonstrate that ExeFuse effectively bridges domain barriers to achieve superior fusion performance. In the future, we plan to enhance the framework by integrating LLMs to further improve the interpretability and reasoning depth of the neuro-symbolic execution process.